An unknown gas is found to be 2.25 times faster than Diphosphorus Pentoxide. What is the molar mass of the unknown gas?
Use four significant figures AND do NOT include units
1
Expert's answer
2020-04-28T12:21:29-0400

R1/R2=√M2/M1

R1/R2=2.25

2.25=√M2/M1

M1=M2/5.0625

M1= 142/5.0625 = 28.